A uniform magnetic field its intensity is 3x10-3 T passes an area of the coil with 100 turns which makes 60 angles with the coil. What is the area of the coil if the magnetic flux is 0.15 Wb?


1
Expert's answer
2022-05-15T17:02:48-0400
"\Phi=NBA\cos\theta"

"A=\frac{\Phi}{NB\cos\theta}"

"A=\frac{0.15}{100*3*10^{-3}\cos30^\circ}=0.577\:\rm m^2"